a pharmacist has 700ml of a solution that is 70% active ingredient. how much pure water should she add to the solution to get a solution that is 10% active ingredient?
Found 2 solutions by TimothyLamb, ewatrrr:
Answer by TimothyLamb(4379)

x = ml of pure water
y = ml of 70% ingredient/water = ml of 30% water/ingredient
z = ml of 10% ingredient/water = ml of 90% water/ingredient
---
y = 700
1x + 0.30y = 0.90z
z = x + y
---
put the system of linear equations into standard form
---
y = 700
1x + 0.30y - 0.90z = 0
x + y - z = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = ml of pure water = 4200
y = 700
z = 4900
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---

Answer by ewatrrr(24785)

x amount of pure water added
water = water
...
x + .30(700ml) = .90(700ml + x)
x = .60(700ml)/.10 = 4200ml, amount of pure water added